31. Oktober 2017: Dienstag

Today’s Activities:

  • Backups:
    • bujbod1 had an NFS outage that was coincidental with a smartd check
      at 4pm on Monday – look into adjusting smartd parameters.
    • SR 09053113 (Sev 2) from last Friday
      • EMC requested information, telling me that it
        is a known bug and is fixed in 8.2.4.9 (which is what I am using)
        So, obviously, not fixed.
    • Reclaimed three (3) tapes for the vaults.
    • TAPE clones completed at 18:15 (yesterday).
    • DISK clones completed at 16:21 (yesterday).
    • Another tweak to the staging script.
      Besides not logging that we ran out of savesets to move,
      do not try to move zero save sets.
  • LDAP/Accounts:
    • Issues from last night’s update,
      one is left over from yesterday,
      other is another possible former employee come back with a new id.
      Waiting on HR response to both.
    • Cleaned up an email routing issue with IAM.

Tags:

Leave a Reply

You must be logged in to post a comment.